OpenStudy (anonymous):

Write an expression for the general or nth term, an, for the geometric sequence. 1, -4, 16, -64... A. An=1.(-4)^n-1 B. An=1.(-4)^n C. An=1.4^n-1 D. An=1.4^n

OpenStudy (freckles):

so have you notice the common ratio for example what are the following equal to: \[\frac{-4}{1}=? \\ \frac{16}{-4}=? \\ \frac{-64}{16}=?\] notice I just did term/previous term

OpenStudy (anonymous):

Is it 4?

OpenStudy (freckles):

you should find that those all equal to -4 which they should have term/previous term hold a constant number since they claimed it was a geometric sequence but anyways since the common ratio is -4 you can definitely rule out C and D and only look at A and B

OpenStudy (freckles):

Now the only way I can answer if it is A or B is if you tell me if your class uses n=0 for the first term or you use n=1 for the first term

OpenStudy (freckles):

I bet it is n=1

OpenStudy (freckles):

but if you know this will really be needed because both of those could described the sequence depending on how n is defined

OpenStudy (freckles):

anyways if it is n=1 which of the following gives 1 as the answer: \[A_1=1(-4)^{1-1} \text{ using choice A here } \\ A_1=1(-4)^{1} \text{ using choice B here }\]

OpenStudy (freckles):

we are looking to see which gives A_1 as 1 since 1 is the first number in the sequence

OpenStudy (anonymous):

Okay, thanks. I think it's A.

OpenStudy (freckles):

yes that is only if your class uses n=1 to give first term